An internal control policy for an online fashion accessories store requires a quality assurance check before a shipment is made. The tolerable exception rate for this internal control is 0.05. During a n audit, 500 shipping records were sampled from a population of 5,000 shipping records and 12 were found that violated the control. What is the upper bound for a 95% one-sided confidence interval estimate for the rate of noncompliance?
Could you please help me?? Thanks Rivers

Answer by stanbon(75887)

An internal control policy for an online fashion accessories store requires a quality assurance check before a shipment is made. The tolerable exception rate for this internal control is 0.05. During an audit, 500 shipping records were sampled from a population of 5,000 shipping records and 12 were found that violated the control. What is the upper bound for a 95% one-sided confidence interval estimate for the rate of noncompliance?
---------------
The sample proportion: p'= 12/500 = 0.024
The standard error: E = 1.645*sqrt[0.024*0.076/500] = 0.011...
-------------------
Upper bound of the 95% confidence interval: p'+E = 0.024+0.011 = 0.035
========================
Cheers,
Stan H.